# The exterior derivative and the mean value equality in $\mathbb{R}^{n}$

### v2

@@ 1 Introduction
In this survey article, we revisit a perspective on the exterior derivative operator on differential forms that yields a higher-dimensional generalisation of the Mean Value Theorem (MVT) as an equality, in contrast with the traditional presentation in vector calculus texts, where the theorem typically appears only in the form of an inequality. These ideas were first formulated by Felipe Acker in his somewhat unknown articles [1, 2]. We propose here a thorough revision of that theory, clarifying some imprecisions in the original proofs and using it to deduce Stokes’ theorem via more accessible and elementary arguments. While stronger formulations do exist – most notably within the framework of modern Geometric Measure Theory in terms of normal currents, cf. [3, §4], our aim is to present these results in an easily approachable language.
@@
The conceptual key is a geometric interpretation of the exterior derivative as an ‘infinitesimal flux’ (Definition 11), much in the vein of the ‘physical’ interpretation of the divergence of a vector field (see, e.g., [4, p. 189]). This alternative definition allows e.g. for the rigorous ‘differentiation’ of certain discontinuous differential forms, and it coincides with the usual notion of exterior derivative in the differentiable case (Theorem 8). As an illustration, consider the following lemma for 0-forms in dimension $1$ (the proof is elementary and can be found at the end of §3.4):
@@
In the general setting of $k$-forms in dimension $n$, we mimic Lemma 2 and define the exterior derivative $\text{\rm D}\omega_{x}$ of $\omega$ at the point $x$ by requiring that, for every choice of directions $v_{1},\dots,v_{k}\in\mathbb{R}^{n}$, the value of $(\text{\rm D}\omega)_{x}(v_{1},\dots,v_{k})$ be given by the limit of the normalised flux of $\omega$ across the boundary of $C^{1}$-parametrised $k$-blocks shrinking to $x$ (see Definition 11). In dimensions $n>1$ and degree $k>0$, the right-hand side of this definition still makes sense even when $\omega$ is not continuous at $x$, let alone differentiable (see e.g. Example 2). Moreover, we prove that in the differentiable case this flux-based exterior derivative $\text{\rm D}\omega$ agrees with the usual exterior derivative $\text{\rm d}\omega$ (see Theorem 8). Thus, the definition extends the classical operator d to a strictly larger class of differential forms. Moreover, with this viewpoint, we can prove the following generalisation of MVT as an equality (see Theorem 5):
@@ 2 Repercussions of the Trisection Lemma
We will discuss a rather non-standard proof (cf. [5, p. 160] and [1, p. 210]) of the one-dimensional Mean Value Theorem. The essence of all generalisations we carry out in this survey lies in the Trisection Lemma stated in the Introduction (Lemma 1), so let us begin with its proof.
@@ 2.1 The mean value theorem and the fundamental theorem of calculus
Let us recall the statement of the one-dimensional version of the MVT (outlined in [5, p. 160] and [1, p. 210]) and its proof using the Trisection Lemma:
@@ 2.2 Chains, blocks and integration in $\mathbb{R}^{n}$
An important property of the boundary operation $\partial$ is that $\partial(\partial c)=0$, for every $k$-chain $c$ on some open subset $A$ of $\mathbb{R}^{n}$. The interested reader can consult [6, Theorem 4-12, p.99] for a proof of this fact.
@@ 2.3 Generalisation of the Trisection Lemma
Now we know in what sense to expect a generalisation of the MVT, the first concrete task is to extend Lemma 1 (see [5, p. 274] and [2, p.5] for more informal proofs, and [1, p. 213] for a two-dimensional version).
@@
Note in the above proof that we only needed the continuity of $\omega$ to ensure the continuity of $m$. Therefore, the relevant continuity is not that of $\omega$ in the usual sense, but a weaker one concerning the pairing between differential forms and chains given by the integral. Given a (k1)(k-1)-form $\omega$, we want to ensure that the integral of $\omega$ on the boundary of nearby $k$-blocks takes nearby values. This motivates the following definition (see [5, p. 273]; also compare with [1, p. 213] and [2, p.4]).
@@
Thus, what we have actually proved in the proof of Lemma 6 is the following (cf. [5, p. 274]):
@@ 3 The exterior derivative as an infinitesimal flux
We recall some standard definitions to fix notation. Let $\omega\in\Omega^{k}(U)$, where $U\subset\mathbb{R}^{n}$ is an open subset, and write
@@
where the sum is taken over all increasing multi-indices $I\coloneqq\{i_{1}<\ldots<i_{k}\}\subseteq\{1,\ldots,n\}$, we have $\text{\rm d}x^{I}:=\text{\rm d}x^{i_{1}}\wedge\ldots\wedge\text{\rm d}x^{i_{k}}$, and the component functions $\omega_{I}:U\rightarrow\mathbb{R}$ are given by
$\omega_{I}(x)=\omega_{x}(\partial_{i_{1}},\ldots,\partial_{i_{k}}).$
@@
$\omega^{\prime}(x):\mathbb{R}^{n}\rightarrow\Lambda^{k}(\mathbb{R}^{n})^{\ast},$
@@
$|\eta(v_{1},\ldots,v_{k})|\leq\left\|\eta\right\||v_{1}|\ldots|v_{k}|,\quad\forall v_{1},\ldots,v_{k}\in\mathbb{R}^{n}.$ (4)
@@ 3.1 The standard definition of the exterior derivative
Let $\omega$ be a $k$-form on some open subset $U$ of $\mathbb{R}^{n}$. If $\omega$ is differentiable at some point $x\in U$, then we define the (usual) exterior derivative of $\omega$ at $x$ to be the $(k+1)$-linear form $\text{\rm d}\omega_{x}$ in $\mathbb{R}^{n}$ given by
@@
The equivalence is straightforward, writing $v_{i}=\sum_{j}v_{i}^{j}e_{j}$ in the canonical basis, for each $i=1,\ldots,k$}, and noting that:
@@
    $\omega^{\prime}(x)v_{i}=\sum_{j}v_{i}^{j}\omega^{\prime}(x)e_{j}=\sum_{I,j}v_{i}^{j}\frac{\partial\omega_{I}}{\partial x^{j}}{\text{\rm d}}x^{I};$
  • for each multi-index $I=\{i_{1}<\ldots<i_{k}\}$ and $j=1,\ldots,k$, we have
@@ 3.2 The infinitesimal flux interpretation
Before presenting Acker’s definition of the exterior derivative (cf. [5, p. 279]), we first provide a heuristic argument, motivated by attempting to reverse-engineer the property we wish to obtain. The classical Stokes’ theorem (see [6, Theorem 4-13, p. 102]) for a differential (k1)(k-1)-form $\omega$ of class $C^{1}$ on an open set $U\subset\mathbb{R}^{n}$ asserts that, for every $C^{1}$ singular $k$-block $c:B^{k}\to U$, one has
@@
  • The strange condition $L(B)/l(B)<K$ implies that the ratio $\text{diam}(B)\cdot\text{area}(\partial B)/\text{vol}(B)$ remains bounded above; this will play an important role in showing that the above definition of the exterior derivative coincides with the usual one in the differentiable case (Theorem 8).
  • This definition is very geometric and intuitive: the exterior derivative of $\omega$ at a point $x$ is the limit of the “average” flux of $\omega$ across the boundary of small blocks shrinking (nicely) to $x$. This contrasts with Definition 9, which involves a somewhat computational notion of differentiability. Here lies an important point: the distinction between the notions of differentiability and derivability.
  • The definition does not even require $\omega$ to be continuous at $x$ to ensure the existence of the exterior derivative: indeed, the $1$-form $\omega$ defined on $\mathbb{R}^{2}$ by $\omega=f\text{\rm d}x^{1}$, where
@@
    is discontinuous at the origin but has a well-defined exterior derivative $\text{\rm D}\omega_{(x_{1},x_{2})}=0$ at any point $x=(x_{1},x_{2})\in\mathbb{R}^{2}$; see Example 2.
@@
Fix any point $x\in U$ and a $C^{1}$ map $\varphi:A\to U$ with $\varphi(p)=x$, and let $B\subset A$ be a $k$-block containing $p$. Denote by $c_{ij}$ the singular (k1)(k-1)-faces of $B$ as in Definition 4. Then, for each face $c_{ij}$, the pullback $c_{ij}^{*}\omega$ is constant on the face except possibly on the preimage $c_{ij}^{-1}(S)$. Now, under our assumptions, by the argument in Step 3 of Remark 4, the integral of the pullback over $c_{ij}^{-1}(S)$ vanishes. Moreover, as for the constant part of $c_{ij}^{*}\omega$, it gives equal and opposite contributions on opposite faces due to the alternating signs in the boundary sum:
@@
and taking the limit $B\to p$ in Definition 11 gives
@@ 3.3 The mean value equality for differential forms
The results discussed in the remainder of Section 3 stem from the bibliography according to a pattern: they are stated and argued for in [5, 2] and proved as a low-dimensional instance in [1]. We therefore propose a systematic exposition of the theory with self-contained and fully rigorous proofs of the general statements. Let us begin with the mean value theorem for differential forms:
@@
One can cast these last results in a more general setting, given the good behavior of D under pullbacks (claimed without proof in [5, p. 286]):
@@
The naturality of D makes rigorous the following generalised version of Stokes’ theorem, claimed in [5, p. 281, Theorem A] (see also [2, p.7] for a manifold version):
@@ 4 Stokes’ theorem
We can easily derive some other well-known results. In what follows, by a (closed) rectangle in $\mathbb{R}^{2}$ we mean a possibly rotated and translated $2$-block, i.e. the image of a $2$-singular block $c:B\to\mathbb{R}^{2}$ of the form $c=E\circ i$, where $i:B\hookrightarrow\mathbb{R}^{2}$ is the inclusion of the $2$-block $B$ and $E:\mathbb{R}^{2}\to\mathbb{R}^{2}$ is a rigid motion in $\mathbb{R}^{2}$ (i.e. the composition of a translation with a rotation on the plane). In particular, we obtain (see [1, p. 215]) the following version of Green’s theorem (on rectangles):
@@ 5 Numerical Exterior Derivative
Building on the exposition in Section 3, we propose a numerical method for computing the components of the exterior derivative $\text{\rm D}\omega$ of a given (k1)(k-1)-form $\omega$ in an open set $U\subset\mathbb{R}^{n}$. Our method operates at any point $x\in U\subset\mathbb{R}^{n}$ using a black-box numerical sampler that does not require access to the analytical expressions of the components of $\omega$. In addition, a demo implementation in SageMath [7] for exterior differentiation in $\mathbb{R}^{3}$ is available at github.com/TomasSilva/NumericalExteriorDerivative.
Contrary to existing approaches, such as Discrete Exterior Calculus (DEC) [8], which computes the exterior derivative via discretization over a mesh, or Finite Element Exterior Calculus (FEEC) [9], that uses the structure of the discrete de Rham complex to approximate differentiation, our method treats $\omega$ as a black-box function and does not require any mesh decomposition. Indeed, we assume access only to numerical evaluations of $\omega$ as a skew-symmetric (0,k1)(0,k-1)-tensor represented by an $n\times\cdots\times n$ array with k1k-1 skew-symmetric slots at any given point in $\mathbb{R}^{n}$.
@@
Acknowledgements The main ideas developed in §§24 were originally proposed by Felipe Acker in [5, 1, 2], where they are proved in somewhat flexible terms; above all, the most comprehensive source [5] is only available in Portuguese. So we dedicate this article to Prof. Acker and hope to bring his innovation, now in a fully rigorous form, to a much broader audience.

### v3

@@ 1 Introduction
In this survey article, we revisit a perspective on the exterior derivative operator on differential forms that yields a higher-dimensional generalisation of the Mean Value Theorem (MVT) as an equality, in contrast with the traditional presentation in vector calculus texts, where the theorem typically appears only in the form of an inequality. These ideas were first formulated by Felipe Acker in his somewhat unknown articles [paper, paper2]. We propose here a thorough revision of that theory, clarifying some imprecisions in the original proofs and using it to deduce Stokes’ theorem via more accessible and elementary arguments. While stronger formulations do exist – most notably within the framework of modern Geometric Measure Theory in terms of normal currents, cf. [Federer1996, §4], our aim is to present these results in an easily approachable language.
@@
The conceptual key is a geometric interpretation of the exterior derivative as an ‘infinitesimal flux’ (Definition 11), much in the vein of the ‘physical’ interpretation of the divergence of a vector field (see, e.g., [Arnold1989, p. 189]). This alternative definition allows e.g. for the rigorous ‘differentiation’ of certain discontinuous differential forms, and it coincides with the usual notion of exterior derivative in the differentiable case (Theorem 8). As an illustration, consider the following lemma for 0-forms in dimension $1$ (the proof is elementary and can be found at the end of §3.4):
@@
In the general setting of (k1)(k-1)-forms in dimension $n$, we mimic Lemma 2 and define the exterior derivative $\text{\rm D}\omega_{x}$ of $\omega$ at the point $x$ by requiring that, for every choice of directions $v_{1},\dots,v_{k}\in\mathbb{R}^{n}$, the value of $(\text{\rm D}\omega)_{x}(v_{1},\dots,v_{k})$ be given by the limit of the normalised flux of $\omega$ across the boundary of $C^{1}$-parametrised $k$-blocks shrinking to $x$ (see Definition 11). In dimensions $n>1$ and degree $k>1$, the right-hand side of this definition still makes sense even when $\omega$ is not continuous at $x$, let alone differentiable (see e.g. Example 2). Moreover, we prove that in the differentiable case this flux-based exterior derivative $\text{\rm D}\omega$ agrees with the usual exterior derivative $\text{\rm d}\omega$ (see Theorem 8). Thus, the definition extends the classical operator d to a strictly larger class of differential forms. Moreover, with this viewpoint, we can prove the following generalisation of MVT as an equality (see Theorem 5):
@@
Organisation of the paper. Section 2 develops the subdivision framework on singular chains that makes it possible to extend to higher dimensions the one-dimensional subdivision argument underlying the Fundamental Theorem of Calculus. The Trisection Lemma and its higher-dimensional extensions provide a systematic decomposition of chains into controlled pieces, yielding in particular the notion of flux-continuity. These results furnish the geometric control on boundary integrals that is essential for the limiting arguments used throughout the paper.
Section 3 introduces the exterior derivative as an infinitesimal flux. Following the paradigm outlined in the introduction, we define the operator D via normalised boundary integrals over shrinking blocks and establish its fundamental properties, including a mean value equality for differential forms, a Stokes formula for D, the identity $\text{\rm D}^{2}=0$, and compatibility with the classical exterior derivative in the differentiable setting.
Section 4 combines these results to obtain a Stokes theorem for the classical exterior derivative under hypotheses directly analogous to those of the Fundamental Theorem of Calculus.
Finally, Section 5 presents a numerical application of the theory. The flux interpretation leads to a natural and easily implementable approximation scheme for the exterior derivative based on boundary integrals, illustrating that this viewpoint is not only conceptually clarifying but also computationally effective.
@@ 2 Repercussions of the Trisection Lemma
We will discuss a rather non-standard proof (cf. [acker1, p. 160] and [paper, p. 210]) of the one-dimensional Mean Value Theorem. The essence of all generalisations we carry out in this survey lies in the Trisection Lemma stated in the Introduction (Lemma 1), so let us begin with its proof.
@@ 2.1 The mean value theorem and the fundamental theorem of calculus
Let us recall the statement of the one-dimensional version of the MVT (outlined in [acker1, p. 160] and [paper, p. 210]) and its proof using the Trisection Lemma:
@@ 2.2 Chains, blocks and integration in $\mathbb{R}^{n}$
An important property of the boundary operation $\partial$ is that $\partial(\partial c)=0$, for every $k$-chain $c$ on some open subset $A$ of $\mathbb{R}^{n}$. The interested reader can consult [spivak, Theorem 4-12, p.99] for a proof of this fact.
@@ 2.3 Generalisation of the Trisection Lemma
Now we know in what sense to expect a generalisation of the MVT, the first concrete task is to extend Lemma 1 (see [acker1, p. 274] and [paper2, p.5] for more informal proofs, and [paper, p. 213] for a two-dimensional version).
@@
Note in the above proof that we only needed the continuity of $\omega$ to ensure the continuity of $m$. Therefore, the relevant continuity is not that of $\omega$ in the usual sense, but a weaker one concerning the pairing between differential forms and chains given by the integral. Given a (k1)(k-1)-form $\omega$, we want to ensure that the integral of $\omega$ on the boundary of nearby $k$-blocks takes nearby values. This motivates the following definition (see [acker1, p. 273]; also compare with [paper, p. 213] and [paper2, p.4]).
@@
Thus, what we have actually proved in the proof of Lemma 6 is the following (cf. [acker1, p. 274]):
@@ 3 The exterior derivative as an infinitesimal flux
We recall some standard definitions to fix notation. Let $\omega\in\Omega^{k-1}(U)$, where $U\subset\mathbb{R}^{n}$ is an open subset, and write
@@
where the sum is taken over all increasing multi-indices $I\coloneqq\{i_{1}<\ldots<i_{k-1}\}\subseteq\{1,\ldots,n\}$, we have $\text{\rm d}x^{I}:=\text{\rm d}x^{i_{1}}\wedge\ldots\wedge\text{\rm d}x^{i_{k-1}}$, and the component functions $\omega_{I}:U\rightarrow\mathbb{R}$ are given by
$\omega_{I}(x)=\omega_{x}(\partial_{i_{1}},\ldots,\partial_{i_{k-1}}).$
@@
$\omega^{\prime}(x):\mathbb{R}^{n}\rightarrow\Lambda^{k-1}(\mathbb{R}^{n})^{\ast},$
@@
$|\eta(v_{1},\ldots,v_{k-1})|\leq\left\|\eta\right\||v_{1}|\ldots|v_{k-1}|,\quad\forall v_{1},\ldots,v_{k-1}\in\mathbb{R}^{n}.$ (4)
@@ 3.1 The standard definition of the exterior derivative
Let $\omega$ be a (k1)(k-1)-form on some open subset $U$ of $\mathbb{R}^{n}$. If $\omega$ is differentiable at some point $x\in U$, then we define the (usual) exterior derivative of $\omega$ at $x$ to be the $k$-linear form $\text{\rm d}\omega_{x}$ in $\mathbb{R}^{n}$ given by
@@
The equivalence is straightforward, writing $v_{i}=\sum_{j=1}^{n}v_{i}^{j}e_{j}$ in the canonical basis, for each $i=1,\ldots,k$}, and noting that:
@@
    $\omega^{\prime}(x)v_{i}=\sum_{j=1}^{n}v_{i}^{j}\omega^{\prime}(x)e_{j}=\sum_{I,j}v_{i}^{j}\frac{\partial\omega_{I}}{\partial x^{j}}{\text{\rm d}}x^{I};$
  • for each multi-index $I=\{i_{1}<\ldots<i_{k-1}\}$ and $j=1,\ldots,n$, we have
@@ 3.2 The infinitesimal flux interpretation
Before presenting Acker’s definition of the exterior derivative (cf. [acker1, p. 279]), we first provide a heuristic argument, motivated by attempting to reverse-engineer the property we wish to obtain. The classical Stokes’ theorem (see [spivak, Theorem 4-13, p. 102]) for a differential (k1)(k-1)-form $\omega$ of class $C^{1}$ on an open set $U\subset\mathbb{R}^{n}$ asserts that, for every $C^{1}$ singular $k$-block $c:B^{k}\to U$, one has
@@
  • Although Definition 11 requires the limit to be taken uniformly over all $C^{1}$ parametrisations $\varphi$ with prescribed first-order data at $p$, this condition is natural from the geometric viewpoint: it ensures that the resulting operator is compatible with arbitrary singular chains and depends only on the first-order behaviour of $\varphi$ at $p$.
  • The condition $L(B)/l(B)<K$ ensures a uniform bound on the ratio $\text{diam}(B)\cdot\text{area}(\partial B)/\text{vol}(B)$; this will play an important role in showing that the above definition of the exterior derivative coincides with the usual one in the differentiable case (Theorem 8).
  • This definition is very geometric: the exterior derivative of $\omega$ at a point $x$ is the limit of the “average” flux of $\omega$ across the boundary of small blocks shrinking (in a controlled way) to $x$. This contrasts with Definition 9, which involves a somewhat computational notion of differentiability. Here lies an important point: the distinction between the notions of differentiability and derivability.
  • The notion of derivability is strictly weaker than classical differentiability. In particular, $\omega$ need not be continuous at $x$ for $\text{\rm D}\omega_{x}$ to exist. For instance, the $1$-form $\omega$ defined on $\mathbb{R}^{2}$ by $\omega=f\text{\rm d}x^{1}$, where
@@
    is discontinuous at the origin but has a well-defined exterior derivative $\text{\rm D}\omega_{(x_{1},x_{2})}=0$ at every point $x=(x_{1},x_{2})\in\mathbb{R}^{2}$; see Example 2.
@@
Fix any point $x\in U$ and a $C^{1}$ map $\varphi:A\to U$ with $\varphi(p)=x$, and let $B\subset A$ be a $k$-block containing $p$. Denote by $c_{ij}$ the singular (k1)(k-1)-faces of $B$ as in Definition 4. Then, for each face $c_{ij}$, the pullback $c_{ij}^{*}\omega$ is constant on the face, except possibly on the preimage $c_{ij}^{-1}(S)$. Now, under our assumptions, by the argument in Step 3 of Remark 4, the integral of the pullback over $c_{ij}^{-1}(S)$ vanishes. Moreover, as for the constant part of $c_{ij}^{*}\omega$, it gives equal and opposite contributions on opposite faces due to the alternating signs in the boundary sum:
@@
and taking the limit as $B\to p$ in Definition 11 yields
@@ 3.3 The mean value equality for differential forms
The results discussed in the remainder of Section 3 stem from the bibliography according to a pattern: they are stated and argued for in [acker1, paper2] and proved as a low-dimensional instance in [paper]. We therefore propose a systematic exposition of the theory with self-contained and fully rigorous proofs of the general statements. Let us begin with the mean value theorem for differential forms:
@@
One can cast these last results in a more general setting, given the good behavior of D under pullbacks (claimed without proof in [acker1, p. 286]):
@@
The naturality of D makes rigorous the following generalised version of Stokes’ theorem, claimed in [acker1, p. 281, Theorem A] (see also [paper2, p.7] for a manifold version):
@@ 4 Stokes’ theorem
We can easily derive some other well-known results. In what follows, by a (closed) rectangle in $\mathbb{R}^{2}$ we mean a possibly rotated and translated $2$-block, i.e. the image of a $2$-singular block $c:B\to\mathbb{R}^{2}$ of the form $c=E\circ i$, where $i:B\hookrightarrow\mathbb{R}^{2}$ is the inclusion of the $2$-block $B$ and $E:\mathbb{R}^{2}\to\mathbb{R}^{2}$ is a rigid motion in $\mathbb{R}^{2}$ (i.e. the composition of a translation with a rotation on the plane). In particular, we obtain (see [paper, p. 215]) the following version of Green’s theorem (on rectangles):
@@ 5 Numerical Exterior Derivative
Building on the exposition in Section 3, we propose a numerical method for computing the components of the exterior derivative $\text{\rm D}\omega$ of a given (k1)(k-1)-form $\omega$ in an open set $U\subset\mathbb{R}^{n}$. Our method operates at any point $x\in U\subset\mathbb{R}^{n}$ using a black-box numerical sampler that does not require access to the analytical expressions of the components of $\omega$. In addition, a demo implementation in SageMath [sagemath] for exterior differentiation in $\mathbb{R}^{3}$ is available at github.com/TomasSilva/NumericalExteriorDerivative.
Contrary to existing approaches, such as Discrete Exterior Calculus (DEC) [Hirani], which computes the exterior derivative via discretization over a mesh, or Finite Element Exterior Calculus (FEEC) [FEEC], that uses the structure of the discrete de Rham complex to approximate differentiation, our method treats $\omega$ as a black-box function and does not require any mesh decomposition. Indeed, we assume access only to numerical evaluations of $\omega$ as a skew-symmetric (0,k1)(0,k-1)-tensor represented by an $n\times\cdots\times n$ array with k1k-1 skew-symmetric slots at any given point in $\mathbb{R}^{n}$.
@@
Acknowledgements The main ideas developed in §§24 were originally proposed by Felipe Acker in [acker1, paper, paper2], where they are proved in somewhat flexible terms; above all, the most comprehensive source [acker1] is only available in Portuguese. So we dedicate this article to Prof. Acker and hope to bring his innovation, now in a fully rigorous form, to a much broader audience.